# Exploring Kolmogorov-Arnold Networks for Interpretable Time Series Classification

### v2

@@ Abstract
Time series classification is a relevant step supporting decision-making processes in various domains, and deep neural models have shown promising performance. Despite significant advancements in deep learning, the theoretical understanding of how and why complex architectures function remains limited, prompting the need for more interpretable models. Recently, the Kolmogorov-Arnold Networks (KANs) have been proposed as a more interpretable alternative. While KAN-related research is significantly rising, to date, the study of KAN architectures for time series classification has been limited. In this paper, we aim to conduct a comprehensive and robust exploration of the KAN architecture for time series classification on the UCR benchmark. More specifically, we look at a) how reference architectures for forecasting transfer to classification, at the b) hyperparameter and implementation influence on the classification performance in view of finding the one that generalizes best on the classification benchmark, the c) complexity trade-offs and d) interpretability advantages. Our results show that (1) Efficient KAN outperforms MLP in performance and computational efficiency, showcasing its suitability for tasks classification tasks. (2) Efficient KAN is more stable than KAN across grid sizes, depths, and layer configurations, particularly with lower learning rates. (3) KAN maintains competitive accuracy compared to state-of-the-art models like HIVE-COTE2, with smaller architectures and faster training times, supporting its balance of performance and transparency. (4) The interpretability of the KAN model aligns with findings from SHAP analysis, reinforcing its capacity for transparent decision-making.
Index Terms:
classification, time series, Kolmogorov Arnold networks, multilayer perceptrons
@@ I Introduction
Time series classification is a relevant step supporting decision-making processes in various domains. In healthcare for example, time series classification enables the diagnosis and monitoring of conditions by analyzing patterns in physiological signals such as electrocardiograms or brain activity [1]. In finance, it enhances risk management and fraud detection by classifying trading behaviors and transaction patterns [2]. By classifying activities based on sensor data in human activity recognition, it supports personalized recommendations and safety monitoring [3]. In remote sensing, it contributes to environmental monitoring and land use classification by categorizing satellite or aerial imagery data over time [4]. In these fields, accurate time series classification plays a central role in advancing predictive and analytic capabilities that drive critical outcomes.
Unlike time series forecasting, which aims to predict future data points, time series classification aims to provide accurate interpretation and categorization of temporal data sequences by assigning a categorical label, class label, to each sequence [5]. Following a well-accepted survey [6] that revealed the lack of investigation of deep learning techniques for time series classification, significant progress has been achieved in related research with hundreds of new works, some of them surpassing the non-deep state of the art [7]. According to the most recent and extensive benchmarking work on 100+ datasets from the UCR archive, the most extensive benchmark for time series classification, the state of the art F1 score of 0.886 is achieved by Hive-Cote2 [8] [9], a large ensemble with considerable training time. The non-deep and non-hybrid classification methods that are more explainable at both feature engineering and decision making steps achieve an F1 of up to 0.869 while the deep architectures reach around 0.88 F1 score. Foundation models for time series, based on transformers, are yet to be studied on cross-domain benchmarks such as UCR [10].
@@
The main critique of the original KAN work is concerned with the fairness of the comparison, triggering substantial research into better understanding their overall performance beyond the initial small scale scientific tasks. Subsequently it has been shown that their performance advantage in non-scientific ML tasks such as vision, natural language and audio processing does not hold [16]. Additional work reveals their reduced effectiveness on functions with noise [17] while other works investigate the suitability of replacing the spline functions with wavelets [18]. The performance of KANs vs MLPs on graph learning tasks is compared in [19]. Their preliminary results reveal that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks.
While KAN-related is significantly rising, new scientific works published almost weekly, to date, the study of KAN architectures for time series has been limited. Very early forecasting studies encompass Temporal KANs (TKANs) [20] and Temporal Kolmogorov Arnold Transformer [21] for notional trades, mixture of experts [22] for various domains including traffic and weather, satellite traffic [23] and integrating convolutional layers with KANs to improve time-series forecasting [24]. Furthermore, with respect to time series classification, studies on their robustness are emerging [25].
In this paper we aim to conduct a comprehensive and robust exploration of the KAN architecture for time series classification on the UCR benchmark. More specifically, we look at a) how the existing architectures for forecasting [23] transfer to classification, at the b) hyperparameter and implementation influence on the classification performance in view of finding the one that performs best on the selected benchmark, the c) complexity trade-ffs and d) interpretability advantages. The contributions of this paper are as follows.
  • A study on the suitability of KAN architectures for classification tasks on the UCR dataset. The study first investigates the feasibility of transferring existing architectures designed for forecasting followed by finding the most suitable architecture for classification.
  • A hyperparameter impact analysis of two KAN implementations, analyzing how variations in grid size, network depth, and node configurations impact classification performance. The analysis leads to finding the best performing KAN configuration for the considered problem.
  • The performance and computational complexity comparison of the original KAN, the Efficient KAN implementation, and Multi-Layer Perceptrons (MLPs) on time series classification tasks reveals EfficientKAN’s superior stability across grid sizes depths and layer configurations.
@@ II Related work
In a real-world satellite traffic forecasting task, KANs achieved comparable or superior accuracy to MLPs while using fewer parameters, showcasing their potential in predictive analytics [23]. This application framed forecasting as a supervised learning problem with specific input-output mappings across time steps, using a GEO satellite traffic dataset. The success of KANs in this setting prompted specialized variants, such as Temporal KAN (T-KAN) and Multivariate Temporal KAN (MT-KAN) [26]. T-KAN targets univariate time series, capturing nonlinear relationships with symbolic regression, while MT-KAN model dependencies between multiple variables for improved accuracy in multivariate settings. Further advancements, like Temporal Kolmogorov-Arnold Networks (TKANs), incorporated LSTM-inspired memory layers, excelling in tasks such as cryptocurrency trading volume forecasting. Temporal Kolmogorov-Arnold Transformer (TKAT) added self-attention mechanisms, outperforming conventional transformers in interpretability and precision in multivariate time series forcasting [20, 21]. Signature-Weighted KANs (SigKAN) extended these innovations by integrating path signatures, making them robust for market volume prediction [27]. The Reversible Mixture of KAN Experts (RMoK) model, another KAN-based approach, introduced a mixture-of-experts structure to assign variables to KAN experts, achieving strong performance in time series forecasting tasks by leveraging temporal feature weights to explain data periodicity [22].
To enhance time series classification robustness, hybrid models that combine KAN and MLP architectures were further explored [25]. The study employed Efficient KAN as the primary implementation of KAN rather than the original implementation. Two hybrid configurations, KAN MLP (KAN with an MLP as the final layer) and MLP KAN (MLP with a KAN as the final layer), were tested across the UCR datasets, and both hybrids achieved performance comparable to traditional KAN and MLP models. Notably, MLP KAN demonstrated increased resilience against adversarial attacks, with this robustness attributed to the lower Lipschitz constant of KAN layers, suggesting that combining KAN with MLP structures can strengthen adversarial resistance in neural networks, thus opening possibilities for more secure and reliable models. Rather than focusing on robustness and hybrid architecture, this study aims at understanding in depth the performance and interpretability of KANs for time series classification.
@@ III Problem Statement
In this paper we analyze the potential of KAN for solving univariate time-series classification problems. A univariate time series $X$ consists of a sequence of observations $[x_{1},x_{2},\ldots,x_{T}]$, where each $x_{t},t\in\{1,2,...,T\}$ represents a value observed at time $t$ representing an ordered set of real values [6].
@@
The classification function $f$ can be realized through a multitude of exiting approaches [7], however we investigate the recently proposed Kolomogrow Arnold Network (KAN) [14] promising increased interpretability and potential computational benefits compared to standard Multilayer Perceptron (MLP).
Figure 1: Example of time series data plotted by class, with time steps on the x-axis and observed values $x_{t}$ on the y-axis. Figure depicts SmoothSubspace dataset from the UCR repository.
@@ III-A KAN for time series classification
[…]
@@
The structure of a KAN can be represented as $[n_{1},…,n_{L+1}]$, where $L$ signifies the total number of layers in the KAN. A deeper KAN can be thus formulated through the composition of $L$ layers as:
$Y=\text{KAN}(X)=(\Phi_{L}\circ\Phi_{L-1}\circ\dots\circ\Phi_{1})X$ (2)
Each layer of a KAN is represented by a matrix where each entry is an activation function. If there is a layer with $d_{in}$ nodes and its neighboring layer with $d_{out}$ nodes, the layer can be represented as a $d_{in}\times d_{out}$ matrix of activation functions (Eq. 4).
$\Phi=\{\phi_{q,p}\},\;p=1,2,\dots,d_{in},\;q=1,2,\dots,d_{out}$ (3)
Unlike traditional MLPs, where activation functions are applied at the nodes themselves, KAN places them at the edges between the nodes. KAN employs the SiLU activation function in combination with B-splines to enhance its expressiveness. This setup allows the edges to control the transformations between layers, while the nodes perform simple summation operations.
$\phi(x)=w_{b}silu(x)+w_{s}\,\text{spline}(x)$ (4)
where SiLU is defined as $\text{silu}(x)=x/(1+e^{-x})$. This activation function allows for smooth, non-linear transformations, which helps the network capture complex patterns in time series data more effectively.
B-splines are smooth polynomial functions of degree (order) $k$, which approximate data using control points. The order $k$ controls the smoothness of the spline. Commonly, $k=3$ is used for cubic splines. Each spline operates over a defined grid $G$, which divides the input space into smaller intervals. The grid is determined by a set of grid points that define the segments of the spline. Increasing the number of grid points, increases spline resolution, enabling more fine grained approximation of the underlying univariate function.
A B-spline of order $k$ requires $G+k$ basis functions to represent the spline over the grid. For each input (node in a layer), the evaluation of a B-spline of order $k$ thus involves computing $G+k-1$ basis functions and performing a weighted sum with the corresponding control points.
@@
To better understand the performance of the models evaluated, it is crucial to explain their decision-making processes. Although black-box models achieve high accuracy, they often lack interpretability. KANs offer a shift in the interpretability paradigm by allowing visualization of the model structure along with its learned B-splines as shown in Fig. 2. Interpretability is further enhanced by model pruning that simplifies visual representation by removing less important connections within the model, which were determined through their significance scores. These scores identify critical connections, which highlight the most important features for the decision of the model. This approach provides insights into how KANs process data and make predictions by focusing on the most influential connections within the structure of the model.
@@ III-A1 KAN learnable parameters
$\text{Parameters}=(d_{in}\cdot d_{out})\cdot(G+k+3)+d_{out}$ (6)
Where $d_{in}$ and $d_{out}$ represent the input and output dimensions of the layer. This is consistent with the findings in the [14], which also emphasize the role of learnable control points and weights associated with the spline functions.
@@ III-A2 KAN FLOPs computation
The total FLOPs for a KAN layer come from three parts: the B-spline transformation, the shortcut path, and the merging of the two branches. Using the De Boor-Cox iterative formulation for B-splines, the FLOPs for one KAN layer in the official KAN implementation is given by [16]:
FLOPs=FLOPs of non-linear $\begin{split}\text{FLOPs}=&\ \text{FLOPs of non-linear function}\cdot d_{\text{in}}\\ &+(d_{\text{in}}\cdot d_{\text{out}})\cdot\left[9\cdot k\cdot(G+1.5\cdot k)\right.\\ &\left.+2\cdot G-2.5\cdot k+3\right]\end{split}$ (7)
FLOPs of one forward pass $FP$ through a network with uniform hidden layers of size $d_{in}\cdot d_{out}=M\cdot M$ are calculated as in Eq. 8:
$\begin{split}N_{FP}=&\ \text{FLOPs of non-linear function}\cdot T\\ &+T\cdot M\cdot\big{[}9\cdot k\cdot(G+1.5\cdot k)\\ &\quad\ +2\cdot G-2.5\cdot k+3\big{]}\\ &+(L-2)\cdot\big{(}\text{FLOPs of non-linear function}\cdot M\\ &\quad\ +M^{2}\cdot\big{[}9\cdot k\cdot(G+1.5\cdot k)\\ &\quad\quad\ +2\cdot G-2.5\cdot k+3\big{]}\big{)}\\ &+\text{FLOPs of non-linear function}\cdot M\\ &+M\cdot C\cdot\big{[}9\cdot k\cdot(G+1.5\cdot k)\\ &\quad\ +2\cdot G-2.5\cdot k+3\big{]}\end{split}$ (8)
III-B MLP time series classification
A Multilayer Perceptron (MLP) is a type of feed-forward artificial neural network, used in various machine learning tasks, including time series classification. MLPs are inspired by the universal approximation theorem, which asserts that such networks can approximate any continuous function given sufficient depth (number of layers) and width (number of neurons per layer).
@@ III-B MLP time series classification
Figure 3: A fully connected MLP architecture.
The architecture of a MLP is depicted in Fig. 3. The MLP consists of multiple layers of nodes (neurons): input layer with $T$ neurons, $K$ hidden layers (layers which are not directly connected to the environment are called hidden [28]) and output layer, consisting of $C$ output neurons.
Each neuron in a layer processes the input it receives from the previous layer by calculating a weighted sum of the inputs, adding a bias term, and then passing the result through an activation function. For a neuron with $n$ inputs, its output is given by the Eq.:
$neuron=a\left(\sum_{i=1}^{n}weight_{i}\cdot input_{i}+bias\right)$ (9)
where $a$ represents the activation function. A commonly used activation function in MLPs is the Rectified Linear Unit (ReLU), defined as $\text{ReLU}(z)=\max(0,z)$.
This structure allows MLPs to model complex, non-linear relationships between inputs and outputs. In time series classification, the input to the MLP $X$, is a fixed-length vector representing the time series data. The network learns to map this input vector to a probability distribution across the possible class labels, where each neuron $p_{c},c\in\{1,2,\dots,C\}$ in the final output layer represents the probability of the input belonging to a specific class $y_{c}$.

### v3

@@ Abstract
Time series classification is a relevant step supporting decision-making processes in various domains, and deep neural models have shown promising performance in this respect. Despite significant advancements in deep learning, the theoretical understanding of how and why complex architectures function remains limited, prompting the need for more interpretable models. Recently, the Kolmogorov-Arnold Networks (KANs) have been proposed as a more interpretable alternative to deep learning. While KAN-related research is significantly rising, to date, the study of KAN architectures for time series classification has been limited. In this paper, we aim to conduct a comprehensive and robust exploration of the KAN architecture for time series classification utilising 117 datasets from UCR benchmark archive, from multiple different domains. More specifically, we investigate a) the transferability of reference architectures designed for regression to classification tasks, b) identifying the hyperparameter and implementation configurations for an architecture that best generalizes across 117 datasets, c) the associated complexity trade-offs and d) evaluate KANs interpretability. Our results demonstrate that (1) the Efficient KAN outperforms MLPs in both performance and training times, showcasing its suitability for classification tasks. (2) Efficient KAN exhibits greater stability than the original KAN across grid sizes, depths, and layer configurations, especially when lower learning rates are employed. (3) KAN achieves competitive accuracy compared to state-of-the-art models such as HIVE-COTE2 and InceptionTime, while maintaining smaller architectures and faster training times, highlighting its favorable balance of performance and transparency. (4) The interpretability of the KAN model, as confirmed by SHAP analysis, reinforces its capacity for transparent decision-making.
@@ I Introduction
Time series classification is a relevant step supporting decision-making processes in various domains. For instance, in healthcare time series classification enables the diagnosis and monitoring of conditions by analyzing patterns in physiological signals such as electrocardiograms or brain activity [1]. In finance, it enhances risk management and fraud detection by classifying trading behaviors and transaction patterns [2]. By classifying activities based on sensor data in human activity recognition, it supports personalized recommendations and safety monitoring [3]. In remote sensing, it contributes to environmental monitoring and land use classification by categorizing satellite or aerial imagery data over time [4]. In these fields, accurate time series classification plays a central role in advancing predictive and analytic capabilities that drive critical decisions.
Unlike time series forecasting, which aims to predict future data points, time series classification aims to provide accurate interpretation and categorization of temporal data sequences by assigning a categorical label or a class label to each sequence [5]. Following a well-accepted survey [6] that revealed the lack of investigation of deep learning techniques for time series classification, significant progress has been achieved in related research with hundreds of new works, some of them surpassing the non-deep state of the art [7]. According to the most recent and extensive benchmarking work on more than 100 datasets from the UCR archive, the most extensive benchmark for time series classification, the state of the art F1 score of 0.886 is achieved by Hive-Cote2 [8] [9], a large ensemble with considerable training time. The non-deep and non-hybrid classification methods that are more explainable at both feature engineering and decision making steps achieve F1 of up to 0.869 while the deep architectures reach around 0.88 F1 score. Foundation models for time series, based on transformers, are yet to be studied on cross-domain benchmarks such as UCR [10].
@@
The main critique of the original KAN work is concerned with the fairness of the comparison, triggering substantial research into better understanding of their overall performance beyond the initial small scale scientific tasks. Subsequently it has been shown that their performance advantage in non-scientific ML tasks such as vision, natural language and audio processing does not hold [16]. Additional work reveals their reduced effectiveness on functions with noise [17] while other works investigate the suitability of replacing the spline functions with wavelets [18]. The performance of KANs vs MLPs on graph learning tasks is compared in [19]. Their preliminary results reveal that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks.
While KAN-related research is significantly rising, with new scientific works published almost weekly, the study of KAN architectures for time series has been limited to date. Very early forecasting studies encompass Temporal KANs (TKANs) [20], Temporal Kolmogorov Arnold Transformer [21] and mixture-of-experts [22] for various domains from traffic and weather to satellite traffic [23] and integrating convolutional layers with KANs to improve time-series forecasting [24]. Furthermore, with respect to time series classification, studies on KANs robustness are emerging [25].
In this paper, we aim to conduct a comprehensive and robust exploration of the KAN architecture for time series classification on the UCR benchmark. More specifically, we look at a) how the existing architectures for forecasting [23] transfer to classification, b) the hyperparameter and implementation influence on the classification performance in view of finding the one that performs best on the selected benchmark, c) the complexity trade-offs, and d) interpretability advantages. The contributions of this paper are as follows.
  • A study on the suitability of KAN architectures for classification tasks on the UCR benchmark consisting of 117 datasets. The study first investigates the feasibility of transferring existing architectures designed for regression followed by finding the most suitable architecture for classification.
  • A hyperparameter impact analysis of two KAN implementations, analyzing how variations in grid size, network depth, and node configurations impact classification performance. The analysis leads to finding the best performing KAN configuration that best generalizes across the 117 datasets.
  • The performance and computational complexity comparison of the original KAN, the Efficient KAN implementation, and Multi-Layer Perceptrons (MLPs) on time series classification tasks revealing Efficient KAN’s superior stability across grid sizes, depths, and layer configurations.
@@ II Related work
In a real-world satellite traffic forecasting task, KANs achieved comparable or superior accuracy to MLPs while using fewer parameters, showcasing their potential in predictive analytics [23]. This application framed forecasting as a supervised learning problem with specific input-output mappings across time steps, using a GEO satellite traffic dataset. The success of KANs in this setting prompted specialized variants, such as Temporal KAN (T-KAN) and Multivariate Temporal KAN (MT-KAN) [26]. T-KAN targets univariate time series, capturing nonlinear relationships with symbolic regression, while MT-KAN models dependencies between multiple variables for improved accuracy in multivariate settings. Further advancements, such as Temporal Kolmogorov-Arnold Networks (TKANs), incorporated LSTM-inspired memory layers, excelling in tasks such as cryptocurrency trading volume forecasting. Temporal Kolmogorov-Arnold Transformer (TKAT) added self-attention mechanisms, outperforming conventional transformers in interpretability and precision in multivariate time series forcasting [20, 21]. Signature-Weighted KANs (SigKAN) extended these innovations by integrating path signatures, making them robust for market volume prediction [27]. The Reversible Mixture of KAN Experts (RMoK) model, another KAN-based approach, introduced a mixture-of-experts structure to assign variables to KAN experts, achieving strong performance in time series forecasting tasks by leveraging temporal feature weights to explain data periodicity [22].
Although T‑KAN, MT‑KAN, TKAN, TKAT and SigKAN extend the Kolmogorov–Arnold mapping, they integrate forecasting‑specific modules (e.g. temporal kernels, attention masks) that diverge from the original design. As a result, they are unfit for classification without stripping or redesigning these components and retraining under a classification objective.
To enhance time series classification robustness, hybrid models that combine KAN and MLP architectures were further explored [25]. The study employed Efficient KAN as the primary implementation of KAN rather than the original implementation. Two hybrid configurations, KAN MLP (KAN with an MLP as the final layer) and MLP KAN (MLP with a KAN as the final layer), were tested across the UCR datasets, and both hybrids achieved performance comparable to traditional KAN and MLP models. Notably, MLP KAN demonstrated increased resilience against adversarial attacks. This robustness is attributed to the lower Lipschitz constant of KAN layers, suggesting that combining KAN with MLP structures can strengthen adversarial resistance in neural networks, thus opening possibilities for more secure and reliable models. Rather than focusing on robustness and hybrid architecture, this study aims at understanding in depth the performance and interpretability of KANs for time series classification.
@@ III Problem Statement
In this paper we analyze the potential of KAN for solving univariate time-series classification problems. A univariate time series $X$ consists of a sequence of observations $x_{1},x_{2},\ldots,x_{T}]$, where each $x_{t},t\in\{1,2,...,T\}$ represents a value observed at time $t$ representing an ordered set of real values [6].
@@
The classification function $f$ can be realized through a multitude of exiting approaches [7], however we investigate the recently proposed Kolmogorov Arnold Network (KAN) [14] promising increased interpretability and potential computational benefits compared to standard Multilayer Perceptron (MLP).
Figure 1: Example of time series data plotted by class, with time steps on the x-axis and observed values $x_{t}$ on the y-axis. Figure depicts SmoothSubspace, one of the 117 datasets from the UCR repository.
@@ III-A KAN for time series classification
Each layer of a KAN is represented by a matrix where each entry is an activation function. If there is a layer with $d_{in}$ nodes and its neighboring layer with $d_{out}$ nodes, the layer can be represented as a $d_{in}\times d_{out}$ matrix of activation functions:
The structure of a KAN can be represented as $n_{1},…,n_{L+1}]$, where $L$ signifies the total number of layers in the KAN. A deeper KAN can be thus formulated through the composition of $L$ layers as:
$Y=\text{KAN}(X)=(\Phi_{L}\circ\Phi_{L-1}\circ\dots\circ\Phi_{1})X.$ (2)
$\Phi=\{\phi_{q,p}\},\;p=1,2,\dots,d_{in},\;q=1,2,\dots,d_{out}.$ (3)
Unlike traditional MLPs, where activation functions are applied at the nodes themselves, KAN places them at the edges between the nodes. KAN employs the SiLU activation function in combination with B-splines to enhance its expressiveness, as per Eq. 4. This setup allows the edges to control the transformations between layers, while the nodes perform simple summation operations.
$\phi(x)=w_{b}silu(x)+w_{s}\,\text{spline}(x).$ (4)
@@
SiLU is defined as $\text{silu}(x)=x/(1+e^{-x})$. This activation function allows for smooth, non-linear transformations, which helps the network capture complex patterns in time series data more effectively.
B-splines are smooth polynomial functions of degree (order) $k$, which approximate data using control points. The order $k$ controls the smoothness of the spline. Commonly, $k=3$ is used for cubic splines. Each spline operates over a defined grid $G$, which divides the input space into smaller intervals. The grid is determined by a set of grid points that define the segments of the spline. Increasing the number of grid points increases spline resolution, enabling more fine grained approximation of the underlying univariate function.
A B-spline of order $k$ requires $G+k$ basis functions to represent the spline over the grid. For each input (node in a layer), the evaluation of a B-spline of order $k$ thus involves computing G+k1G+k-1 basis functions and performing a weighted sum with the corresponding control points.
@@
To better understand the performance of the evaluated models, it is crucial to explain their decision-making processes. Although black-box models achieve high accuracy, they often lack interpretability. KANs offer a shift in the interpretability paradigm by allowing visualization of the model structure along with its learned B-splines as shown in Fig. 2. Interpretability is further enhanced by model pruning that simplifies visual representation by removing less important connections within the model, which were determined through their significance scores. These scores identify critical connections, which highlight the most important features for the decision of the model. This approach provides insights into how KANs process data and make predictions by focusing on the most influential connections within the structure of the model.
@@ III-A1 KAN learnable parameters
$\text{Parameters}=(d_{in}\cdot d_{out})\cdot(G+k+3)+d_{out},$ (6)
where $d_{in}$ and $d_{out}$ represent the input and output dimensions of the layer. This is consistent with findings in [14], which also emphasize the role of learnable control points and weights associated with the spline functions.
@@ III-A2 KAN FLOPs computation
The total FLOPs for a KAN layer come from three parts: the B-spline transformation, the shortcut path, and the merging of the two branches. Using the De Boor-Cox iterative formulation for B-splines, the FLOPs for one KAN layer in the original KAN implementation is given by Eq. 7 [16]:
$\begin{split}\text{FLOPs}=&\ \text{FLOPs of non-linear function}\cdot d_{\text{in}}\\ &+(d_{\text{in}}\cdot d_{\text{out}})\cdot\left[9\cdot k\cdot(G+1.5\cdot k)\right.\\ &\left.+2\cdot G-2.5\cdot k+3\right].\end{split}$ (7)
FLOPs of one forward pass FPFP through a network with uniform hidden layers of size $d_{in}\cdot d_{out}=M\cdot M$ are calculated as in Eq. 8[16]:
@@
III-B Comparison of KAN to MLP for time series classification
A Multilayer Perceptron (MLP) is a very well known type of feed-forward neural network widely used for time series classification.
KANs and MLPs share several architectural principles but differ significantly in how they implement non-linearity and function approximation. Both are fully connected neural networks in which each layer’s nodes are densely connected to the next. The input layer in each architecture corresponds to the length of the input time series, and the output layer produces a probability distribution over class labels, allowing for multiclass classification.
However, their treatment of activation differs notably. In MLPs, activation functions are applied at the neurons (i.e. nodes). In KANs, activation functions are applied on the edges between nodes. Regarding activation functions, the two architectures also use different approaches. MLPs commonly use standard functions such as ReLU, defined as $\text{ReLU}(z)=\max(0,z)$. KANs employ the SiLU (Sigmoid Linear Unit) activation in combination with B-splines.